It took Betsy 10 min more to walk home from school than it took him to walk to school from home. If his total walking time from home to school and back again was 54 min, how long did it take him to walk to school?
-please show in an algebraic equation.

Answer by checkley77(12844)

x+x+10=54
2x=54-10
2x=44
x=44/2
x=22 minutes to get to school &
22+10=32 minutes to walk home from school.